OpenStudy (anonymous):

At a high school, 150 students are surveyed about their electives, a. Find the percent of the students surveyed taking each elective. b. The school population is about 2,500. Estimate the number of students taking each elective. Elective: 35 students take ceramics 17 students take typing 68 students take cooking 30 students other

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I am guessing you would take 35/150=%/100 with each elective for the first question but not sure about the second part of the quesion

OpenStudy (anonymous):

@keegq Me either. That is why I am confused

OpenStudy (anonymous):

trying to find percent of a number

OpenStudy (anonymous):

For ceramics use: 35/130 = x/2500 We assume proportionality. Do the same for the other electives.

OpenStudy (anonymous):

basically what I was going to say ^

OpenStudy (anonymous):

If you find the %, then you can also just do % * 2500. Remember that 2% = 0.02 for example.

OpenStudy (anonymous):

Ok thank you for your help!